OpenStudy (anonymous):

Carson's Bakery and Chocolate Shop is experimenting with making chocolate in the shape of different types of sports equipment. Bayside High School's Volleyball team has ordered 100 solid white chocolate volleyballs. What volume of white chocolate is needed to fill this order if the radius of each volleyball is 4 cm? Rounded to the nearest tenth of a cubic centimeter.

OpenStudy (anonymous):

The equation for the volume of a sphere is\[V = \frac{ 4 }{ 3}\pi r^2\] Substitute r=4

OpenStudy (anonymous):

157.7536 is what i got

OpenStudy (anonymous):

@tabby2014

OpenStudy (anonymous):

I got 67.0206 for one volleyball. You then need to multiply that by 100 to get the total volume of chocolate. So that would be 6702.1 cubic cm.

OpenStudy (anonymous):

is this rounded to the nearest tenth

OpenStudy (anonymous):

its wrong ....@tabby2014